OpenStudy (anonymous):

How to evaluate CSC(15deg) without using a calculator? Thanks.

jimthompson5910 (jim_thompson5910):

by using the identities csc(theta) = 1/sin(theta) sin(A - B) = sin(A)*cos(B) - cos(A)*sin(B)

OpenStudy (anonymous):

Sin(40-30) = (Sin(sqrt(2)/2) (Cos(sqrt(3)/2) - Cos(sqrt(2)/2) (Sin(1/2) = sqrt(6) / 4 - sqrt (2) /4 = sqrt(6) - sqrt(2) / 4 Did I do this right? Thank you so much.

OpenStudy (anonymous):

So it would be: ? How do i simplyfy the fraction within the fraction?